The pulley moves up with velocity of 10m/s.Two blocks are tied by a string which passes over pulley. One block having velocity 5m/s. Find velocity of other block.

Grade:12th pass

2 Answers

Vikas TU
6 years ago
Dear Student,
 
Suppose the tension in one string is T
(T+T)/2=Tension that pulley supports
=>(a1+a2)/2=a of the pulley
=>(v1+v2)/2 = v of he pulley
=>(5+v2)/2=10
=>v2=15m/s.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Eyidf
5 years ago
The correct answer is 25m/s
If u don't know the correct answer then plz don't answerg.